OpenStudy (anonymous):

f o g (x) g o f (x) f(x)=1/2x+8 g(x)=4x+10

OpenStudy (amistre64):

just insert one function into the 'x' spot of the other like this: f(g(x)) = 1 1 ------- = ---------- g(x)+8 (4x +10)+8 f(g(x)) = 1/(4x+18)

OpenStudy (amistre64):

g(f(x)) = 4(f(x))+10 = 4(1/(2x+8))+10 = 4/(2x+8) +10 = 2/(x+4) +10 = (2+10x+10)/(x+4) = (10x+12)/(x+4) if i kept track of it